# Point-MoE: Towards Cross-Domain Generalization in 3D Semantic Segmentation via Mixture-of-Experts

### v1

@@ Abstract
While scaling laws have transformed natural language processing and computer vision, 3D point cloud understanding has yet to reach that stage. This can be attributed to both the comparatively smaller scale of 3D datasets, as well as the disparate sources of the data itself. Point clouds are captured by diverse sensors (e.g., depth cameras, LiDAR) across varied domains (e.g., indoor, outdoor), each introducing unique scanning patterns, sampling densities, and semantic biases. Such domain heterogeneity poses a major barrier towards training unified models at scale, especially under the realistic constraint that domain labels are typically inaccessible at inference time. In this work, we propose Point-MoE, a Mixture-of-Experts architecture designed to enable large-scale, cross-domain generalization in 3D perception. We show that standard point cloud backbones degrade significantly in performance when trained on mixed-domain data, whereas Point-MoE with a simple top-$k$ routing strategy can automatically specialize experts, even without access to domain labels. Our experiments demonstrate that Point-MoE not only outperforms strong multi-domain baselines but also generalizes better to unseen domains. This work highlights a scalable path forward for 3D understanding: letting the model discover structure in diverse 3D data, rather than imposing it via manual curation or domain supervision.
@@ 1 Introduction
In both language and vision, we have witnessed remarkable advances driven by two trends: the aggregation of massive heterogeneous datasets [59, 64, 5, 57, 9, 89], and the deployment of models of ever-increasing capacity trained on them in a unified manner [8, 18, 55, 39, 25, 49]. These models succeed not because they were crafted for a specific domain, but because they were trained to learn underlying patterns across a wide variety of datasets. However, 3D point cloud understanding has yet to follow the same trajectory of progress, despite its wide applications in robotics[13, 50, 84], autonomous systems[15, 14, 44], augmented reality[90, 75], and embodied intelligence[77, 71]. Taking the 3D semantic segmentation task as an example, although diverse 3D datasets exist—such as ScanNet [21], SemanticKITTI [6], nuScenes [11], and ASE [3]—each covers only a narrow subset of the real-world diversity in 3D assets. Point clouds are generated from a variety of pipelines: RGB-D sensors [12], terrestrial and airborne LiDAR [33, 51], and multi-view stereo [83]. Each source brings its own characteristics—differing point densities, coverage patterns, reconstruction artifacts, and semantic biases (Fig. 1 left). The result is a 3D ecosystem of rich, yet siloed data sources, where models trained in one silo often fail to generalize across others.
@@
A natural response to address this challenge is to train across such silos. For example, one might naively train the state-of-the-art Point Transformer V3 (PTv3) [78] (Fig. 1a) on mixed-domain 3D datasets. However, such models struggle to reconcile heterogeneous characteristics of the data, even with increased network capacity (Tab. 1). To mitigate this, recent methods introduce domain-aware components. Point Prompt Training (PPT) [80] (Fig. 1b) modifies PTv3 by assigning domain-specific normalization layers. Similarly, One-for-All [74] achieves improved 3D detection performance by employing a lightweight classifier to infer dataset labels, then applying domain-specific adapters. While these methods show notable gains, they rely on access to domain labels during training and exhibit performance degradation on unseen domains. A likely cause is that adapting only the normalization parameters may be insufficient to fully capture the complex variations required for scalable, cross-domain generalization.
@@
In this work, we hypothesize that Mixture-of-Experts (MoE) architectures [35, 38] are well-suited for cross-domain joint training. We introduce Point-MoE, a sparse MoE model built upon PTv3 [78]. As illustrated in Fig. 1c, Point-MoE replaces each feed-forward layer in PTv3 with a MoE layer comprising of multiple expert networks (implemented as feed-forward MLPs) and a router that dynamically selects a sparse subset of experts for each input token. Our design is motivated by three key insights. First, MoE provides a natural inductive bias for handling diverse inputs by encouraging expert specialization and learning token-to-expert routing [17, 82]. This allows a single model to be trained on mixed-domain data without requiring access to domain labels. Second, compared to normalization-based domain adaptation methods like PPT [80] and One-For-All [74], MoE offers a significantly larger and more expressive representation space, enabling better modeling of domain-specific variations. Third, sparse MoE architectures improve computational efficiency during both training and inference, and have emerged as a standard design for scaling large models [41, 28, 26, 58, 22]. We extend these empirical benefits of MoE, previously established in language and vision, to the domain of 3D point cloud understanding, with a particular focus on 3D semantic segmentation.
@@
Figure 1: Overview of cross-domain training architectures. Point clouds exhibit diverse characteristics across domains. (a) Naively training Point Transformer V3 (PTv3) [78] on multi-domain data leads to degraded performance within each domain. (b) Point Prompt Training (PPT) [80] addresses this by adding domain-aware normalization parameters. (c) Our proposed Point-MoE tackles this challenge with Mixture-of-Experts (MoE), enabling dynamic expert specialization across domains. Domain labels can optionally be provided to the router to guide the expert selection.
We conduct extensive experiments across a diverse mixture of 3D datasets, spanning indoor and outdoor scenes, synthetic and real data, and multiple sensor types. To comprehensively assess the effectiveness of Point-MoE, we evaluate it under both domain-agnostic (i.e., without access to domain labels) and domain-aware training regimes. Our results show that Point-MoE consistently adapts well to both settings, achieving strong generalization to unseen domains while outperforming competitive baselines, with improved training (Fig. 2) and inference efficiency (Tab. 1 and 2). We also conduct a thorough ablation study analyzing key design choices, including the number of experts, sparsity level, and the placement of MoE layers (Tab. 3). Remarkably, even without domain supervision, Point-MoE learns to route input data meaningfully and to specialize experts organically—emerging from the data distribution itself (Fig. 3 and 4).
To the best of our knowledge, this work presents the first systematic exploration of the Mixture-of-Experts architecture for 3D point cloud understanding in the context of cross-domain generalization. These findings point toward a scalable path forward for 3D perception: instead of designing separate models for each dataset or domain, we should aim to develop unified systems that can adapt across the full spectrum of 3D data sources. In this regard, Point-MoE emerges as a promising and flexible framework for bridging domain gaps in large-scale 3D semantic segmentation.
@@ 2 Related Work
Deep learning backbones for 3D scene understanding can be broadly categorized by how they process point clouds: projection-based methods [15, 40, 42, 62], voxel-based methods [48, 61, 19, 29], and point-based methods [52, 53, 67, 86]. More recently, transformer-based architectures operating on point clouds have emerged as strong contenders in this space [31, 79, 87, 78]. Among them, Point Transformer V3 (PTv3) [78] represents the current state of the art, introducing a novel space-filling curve based serialization scheme, that transforms the unstructured 3D points into a 1D sequence, enabling efficient transformer-based modeling. A recent series of works have explored scaling such point cloud models to the multi-dataset setting using domain-aware training strategies. Point Prompt Training (PPT)  [80] demonstrates that using domain-specific normalization layers during pretraining improves 3D segmentation performance across multiple datasets. In 3D detection, One-for-All [74] adopts a sparse convolutional network with domain-specific partitions and normalization layers to boost multi-domain performance. Rather than relying on such domain-specific architectural branches [80, 74], we propose a unified sparse expert framework that learns to specialize dynamically across various domains, without requiring explicit dataset labels during training or inference.
Mixture-of-Experts.
Mixture-of-Experts (MoE) [35, 38] aims to increase model capacity without a proportionate increase in compute costs by levereaging conditional computation – large portions of the model are only activated conditioned on certain input samples [60]. In NLP, MoE has enabled scaling up models from hundreds of millions to hundreds of billions (and beyond) parameters: GShard scales to 600 B parameters [41], Switch Transformer to 1.6 T parameters [28], ST-MoE to 260 B parameters [91], and GLaM to 1.2 T parameters [26], all at roughly the same training cost as a dense model one-tenth the size. V-MoE [58] applies this recipe to vision, demonstrating competitive accuracy at 15 B parameters with batch-priority routing. More recently, Mixtral 8x7B [37] and DeepSeek models [22, 23, 24] all employ Mixture-of-Experts for large language model pretraining.
@@
Apart from being a powerful tool for scaling up model capacity, MoE is well-suited to training on data drawn from heterogeneous domains, because the experts naturally tend to self-organize on a per-domain basis. Early work shows gains in multi-source sentiment and POS tagging [30], domain-adaptive language modeling [32], automatic chart understanding [82], and multilingual translation [17]. Mod-Squad [16] introduces a modular mixture-of-experts framework for multi-task learning, optimizing task-expert matching to balance cooperation and specialization across tasks. DAMEX [36] applies this to visual domains, while MoE-LLaVA [45], DeepSeek-VL [47], and DeepSeek-VL 2 [81] report similar benefits in vision-language models. To our knowledge, ours is the first attempt at extending MoE to address multi-dataset training and cross-domain generalization for 3D point clouds.
3 Point-MoE for Cross-Domain Point Cloud Understanding
@@
Task definition. Given a 3D scene represented as a point cloud $\mathcal{P}=\{p_{i}\}_{i=1}^{n}$, where each point $p_{i}\in\mathbb{R}^{3}$ denotes coordinates, semantic segmentation aims to predict a single class label $\hat{y}_{i}\in\mathcal{C}$ to each point $p_{i}$ from a fixed label set $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{m}\}$. Multi-domain data are utilized during training. Let the domain containing $N$ point clouds be $\mathcal{D}=\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left(p_{i},y_{i}}\right)_{i=1}^{n}}\right\}^{N}$. Then the complete training set across $d$ domains can be represented as: $\mathopen{}\mathclose{{}\left\{\mathcal{D}_{k}}\right\}_{k=1}^{d}$. The goal of multi-domain joint training is to learn a unified model over all domains that minimizes prediction error across $\mathopen{}\mathclose{{}\left\{\mathcal{D}_{k}}\right\}$, while also generalizing effectively to domains unseen during training.
We adopt a minimal Mixture-of-Experts design that closely follows the standard architecture prevalent in the recent NLP literature. This simplicity allows us to leverage existing scalable Transformer-based MoE implementations with minimal modification. For the base model, we use Point Transformer V3 (PTv3) [78], a state-of-the-art architecture for point cloud understanding. Below we introduce the architecture of Point-MoE in detail.
Mixture-of-Experts layer. The core of MoE is to route input tokens dynamically to specialized subnetworks, referred to as experts, via a lightweight gating mechanism. Given an input feature vector $\mathbf{x}\in\mathbb{R}^{d}$, each MoE layer contains $N$ experts $\{f_{i}\}_{i=1}^{N}$ and a gating network $G:\mathbb{R}^{d}\rightarrow\mathbb{R}^{N}$. For each input $\mathbf{x}$, the MoE layer output is computed as a weighted sum over the top-$k$ activated experts:
$\mathrm{MoE}(\mathbf{x})=\sum_{i\in\mathcal{S}_{\mathbf{x}}}G_{i}(\mathbf{x})f_{i}(\mathbf{x})$ (1)
where $\mathcal{S}_{\mathbf{x}}$ denotes the set of top-$k$ experts selected for input $\mathbf{x}$, and $G_{i}(\mathbf{x})$ represents the normalized gating weight assigned to expert $i$. In practice, the gating function is typically implemented via a linear projection followed by a sparse softmax, ensuring activation of only a small subset of experts, with $k\ll N$. Although auxiliary losses for balancing expert load are common in prior work [60], we empirically observed them to be unnecessary and therefore excluded them from our final setup.
Integration into PTv3. To tailor the MoE architecture specifically for 3D point cloud data, we integrate the MoE mechanism directly into PTv3 by replacing the shared linear projection layers in PTv3’s attention module with an MoE layer. The original linear projections mapping input features into key, query, and value vectors within the attention mechanism are replaced by MoE layers, while all subsequent computations remain unchanged. This integration allows the MoE experts to specialize naturally in domain-specific transformations of point cloud data. Fig. 1c provides a visual overview of the modified architecture. It is crucial to insert the MoE layer before the post-normalization; otherwise, we observe a significant reduction in performance gains.
@@
Language-guided classification. To bridge label discrepancies across datasets, we project per-point features to a shared language space using CLIP [56] text embeddings, following PPT [80]. This enables supervision via class names, without relying on dataset labels. For instance, the class “pillow” is missing in ScanNet (grouped as “other”) but is explicitly labeled in Structured3D. Our model aligns point features with the CLIP [56] embedding of “pillow,” allowing knowledge to transfer across domains even in the absence of direct supervision.
Domain-aware gating. Our preliminary experiments show that while basic Point-MoE naturally specializes experts implicitly, providing explicit domain information could further guide and accelerate expert specialization. To explore this, we introduce a dataset-aware gating variant termed Domain-Aware Point-MoE. In this variant, we explicitly associate each training domain $\mathcal{D}_{i}$ with a unique, learnable domain embedding vector $d_{i}\in\mathbb{R}^{d}$. Before routed through the gating mechanism, each input feature $x_{t}$ is concatenated with the corresponding domain embedding: $G_{t}(d)=\mathrm{softmax}(W\cdot(x_{t}\oplus d_{i}))$, where $W$ denotes a learnable linear projection, and the concatenation operation explicitly conditions gating decisions on the domain identity. This explicit conditioning encourages faster and more clearly delineated expert specialization aligned with dataset-specific characteristics observed during training.
Domain randomization for robust gating. While domain-aware gating allows experts to specialize based on known domain identities, this approach assumes access to explicit domain labels at inference time. In practical scenarios, such labels are often unavailable—especially when deploying models in the wild, where domain boundaries may be ambiguous or entirely unknown. While a domain detector can be employed to approximate the domain label by mapping unseen domains to the closest known ones, such detectors may introduce errors or fail to capture nuanced domain shifts. This motivates the need for a more robust mechanism that does not rely on explicit domain identification at test time. To address this, we introduce a domain randomization strategy that enables the model to generalize to unseen or mixed-domain inputs without performance degradation on seen domains.
Specifically, we define a learnable generic domain embedding and randomly assign it to 20% of the training points, replacing their true domain-specific embeddings. This forces the gating mechanism to learn domain-agnostic routing behavior, even in the presence of domain ambiguity. The remaining 80% of points retain their original domain embeddings, ensuring that the model still benefits from explicit supervision where available. In line with prior work [70, 17], we find that domain randomization improves generalization to unseen domains while preserving strong performance on those seen during training.
Training objective. We use Mixture-of-Experts (MoE) in a supervised cross-domain setting, where the goal is to train a unified model $f(x;\theta)$ across $n$ domains $\{\mathcal{D}_{i}\}_{i=1}^{n}$, each containing samples $(X,Y)\sim\mathcal{D}_{i}$. The objective is to minimize the average loss across domains:
$\min_{\theta}\sum_{i=1}^{n}\mathbb{E}_{(X,Y)\sim\mathcal{D}_{i}}\;\mathcal{L}(f(X;\theta),Y)$ (2)
We consider two batching strategies during training: (1) Single-domain batching, where each batch is sampled from only one domain $\mathcal{D}_{i}$ at a time. (2) Mixed-domain batching, where each batch contains samples from multiple domains. Our implementation is based on the official open-source code of PTv3 released by Pointcept [20].

### v2

# Point-MoE: Large-Scale Multi-Dataset Training with Mixture-of-Experts for 3D Semantic Segmentation

@@ Abstract
While massively scaling both data and models have become central in NLP and 2D vision, their benefits for 3D point cloud understanding remain limited. We study the initial step of scaling 3D point cloud understanding under a realistic regime: large-scale multi-dataset joint training for 3D semantic segmentation, with no dataset labels available at training or inference time. Point clouds arise from a wide range of sensors (e.g., depth cameras, LiDAR) and scenes (e.g., indoor, outdoor), yielding heterogeneous scanning patterns, sampling densities, and semantic biases; naively mixing such datasets degrades standard models. Therefore, we introduce Point-MoE, a Mixture-of-Experts design that expands model capacity through sparsely activated expert MLPs and a lightweight top-$k$ router, allowing tokens to select specialized experts without requiring dataset supervision. Trained jointly on a diverse mix of indoor and outdoor datasets, and evaluated on seen datasets as well as in zero-shot settings, Point-MoE outperforms prior methods without using dataset labels for either training or inference. This outlines a scalable path for 3D perception: letting the model discover structure in heterogeneous 3D data rather than imposing it via manual curation or dataset-specific heuristics.
@@ 1 Introduction
In both language and vision, we have witnessed remarkable advances driven by two trends: the aggregation of massive, heterogeneous datasets [61, 66, 4, 58, 8, 97], and the deployment of ever-larger models trained on them in a unified manner [7, 16, 57, 36, 21, 51]. These models succeed not because they are crafted for a specific dataset, but because they learn underlying regularities across multiple datasets. However, 3D point cloud understanding has not yet followed the same trajectory, despite its wide applications in robotics [11, 53, 90], autonomous systems [13, 12, 43], augmented reality [6, 77], and embodied intelligence [79, 72]. Taking 3D semantic segmentation as an example, although diverse 3D datasets exist—such as ScanNet [19], SemanticKITTI [5], nuScenes [9], and ASE [2]—each covers only a narrow slice of real-world variation. Point clouds arise from diverse pipelines (RGB-D, LiDAR, multi-view stereo), each with their own distinct densities, artifacts, and semantic biases (Fig. 1, left). The result is a rich yet siloed ecosystem of 3D models, where models trained on one domain often fail to generalize to others.
@@
A natural response is to train across silos, presumably resulting in a single set of model parameters that is effective across the spectrum of datasets. One might, for example, naively train the state-of-the-art Point Transformer V3 (PTv3) [81] (Fig. 1a) on a mixture of datasets from multiple domains, e.g. indoor and outdoor. In practice, however, such models struggle to reconcile the heterogeneity of the data distribution even as capacity scales (Tab. 1). To mitigate this, recent methods have introduced dataset-aware components. Point Prompt Training (PPT) [82] (Fig. 1b) assigns dataset-specific normalization layers, while One-for-All [75] uses a lightweight dataset classifier with dataset-specific adapters. While effective, these methods rely on dataset labels, identifying which specific dataset an input sample originated from, during training and inference. Moreover, adapting only the normalization parameters may be insufficient for large-scale multi-dataset training.
@@
At deployment, point clouds arrive from mixed, unlabeled sources with unknown provenance, i.e. there is no oracle “dataset ID” available during inference. Accordingly, we study a realistic regime: large-scale multi-dataset joint training for 3D semantic segmentation with no dataset labels at inference. We hypothesize that Mixture-of-Experts (MoE) architectures [30, 34] are well-suited to this setting. We introduce Point-MoE, a sparse MoE model built upon PTv3 [81]. As illustrated in Fig. 1c, Point-MoE replaces each attention projection layer in each PTv3 block with a MoE module composed of expert MLPs and a lightweight router that selects a sparse top-$k$ subset of experts per token. Our design is motivated by three insights. First, MoE encourages expert specialization and token-to-expert routing [15, 87], enabling a single model to learn from mixed datasets without access to dataset labels at training or test time. Second, compared to normalization-based adaptations such as PPT [82] and One-for-All [75], MoE offers far more expressive capacity to model dataset-specific variations. Third, sparse MoE architectures maintain computational efficiency during training and inference, and are a standard mechanism for scaling large models in NLP and 2D vision [40, 24, 22, 59, 20]. We extend these benefits, well-established in language and 2D vision, to 3D point cloud understanding, focusing on 3D semantic segmentation.
@@
Figure 1: Overview of multi-datasets training architectures. Point clouds exhibit diverse characteristics across datasets. (a) Naively training Point Transformer V3 (PTv3) [81] on multi-datasets data leads to degraded performance within each domain. (b) Point Prompt Training (PPT) [82] addresses this by adding dataset-aware normalization parameters. (c) Our proposed Point-MoE tackles this challenge with Mixture-of-Experts (MoE), enabling dynamic expert specialization across datasets.
We conduct extensive experiments across a diverse collection of 3D datasets spanning indoor and outdoor scenes, synthetic and real datasets, and multiple sensor modalities. We evaluate on both in-domain and zero-shot datasets, all without dataset labels. To comprehensively assess Point-MoE, we evaluate both in-domain (seen) datasets and out-of-domain (zero-shot) datasets, all without access to dataset labels. Our goal is to demonstrate that Point-MoE not only achieves strong performance on seen datasets but also generalizes robustly to unseen distributions, without relying on dataset-specific augmentation or supervision. Across both regimes, Point-MoE consistently outperforms competitive multi-dataset baselines. We further ablate the number of experts, sparsity level, and MoE placement (Tab. 3), and find that Point-MoE learns coherent routing patterns and exhibits organic expert specialization driven by the underlying data distribution (Fig. 234).
To the best of our knowledge, this work provides the first systematic study of MoE for 3D point cloud understanding under the large-scale multi-dataset training regime. Our contributions are: (1) we focus on multi-dataset 3D semantic segmentation without domain labels, using a single model, and put forth Mixture-of-Experts as a natural solution to this task; (2) we carefully explore the MoE design space and provide extensive ablations that highlight the trade-offs and inform effective configurations; (3) we demonstrate that Point-MoE achieves state-of-the-art performance across seven datasets under this protocol while maintaining inference efficiency; and (4) we analyze token-level routing trajectories and expert usage, revealing encoder–decoder specialization patterns consistent with 3D geometry. Together, these findings point toward a scalable path for 3D perception: rather than building separate models for each dataset or domain, a single unified system can adapt across the spectrum of 3D data sources, leveraging scaling laws for data and compute [68].
@@ 2 Related Work
Point Cloud understanding. Deep learning backbones for 3D scene understanding can be broadly categorized by how they process point clouds: projection-based methods [13, 38, 41, 64], voxel-based methods [49, 63, 17, 25], and point-based methods [54, 55, 69, 93]. More recently, transformer-based architectures operating on point clouds have emerged as strong contenders in this space [27, 80, 94, 81]. Among them, Point Transformer V3 (PTv3) [81] represents the current state of the art, introducing a novel space-filling curve based serialization scheme, that transforms the unstructured 3D points into a 1D sequence, enabling efficient transformer-based modeling. Recent works [32, 96, 39, 82, 83] scale point cloud understanding models to multi-dataset settings and show that large-scale point cloud data can improve performance on specific datasets: Sonata [83] explores multi-dataset self-supervised learning, but does not train across indoor and outdoor datasets. On the other hand, Point Prompt Training (PPT)  [82] demonstrates that using dataset-specific normalization layers during pretraining improves 3D semantic segmentation performance across multiple datasets along with dataset-specific fine-tuning. In 3D detection, One-for-All [75] adopts a sparse convolutional network with dataset-specific partitions and normalization layers to boost multi-datasets performance, and UniDet3D [37] achieves state-of-the-art performance with multi-dataset joint training by merging label spaces across datasets into unified taxonomy and building upon vanilla transformer [71] architecture. Rather than relying on dataset-specific preprocess or architectural branches [82, 75], we propose a unified sparse expert framework that learns to specialize dynamically across various datasets, without requiring explicit dataset labels during training or inference.
@@
Mixture-of-Experts. Mixture-of-Experts (MoE) [30, 34] increases model capacity without proportional compute by activating only a subset of experts per input [62]. This conditional computation has powered trillion-scale NLP models [40, 24, 98, 22, 33, 20, 46, 47] and demonstrated similar efficiency in vision [59]. Beyond scaling, MoE is naturally suited for heterogeneous or multi-datasets data, as experts tend to self-organize by dataset or task. Prior work has shown improvements in multi-source sentiment analysis and POS tagging [26], dataset-adaptive language modeling [28], automatic chart understanding [87], and multilingual translation [15]. Mod-Squad [14] introduced modular expert routing for multi-task learning, balancing cooperation and specialization. Similar trends are observed in vision and vision-language applications: DAMEX [31], MoE-LLaVA [44] and DeepSeek-VL [48, 85]. To our knowledge, this is the first work to extend MoE to multi-dataset training in 3D point clouds. MoE3D [76] recently introduces a MoE formulation for 3D reconstruction.
Domain generalization. Classical domain generalization (DG) trains on one or more labeled source domains and targets zero-shot performance on a specified target domain, typically in sensor- and semantic-ontology-homogeneous outdoor LiDAR settings [35, 60, 92, 84, 88, 86]. We instead study large-scale joint training on heterogeneous datasets with naturally misaligned label spaces, evaluating a single model on both seen and zero-shot datasets. Unlike DG, our goal is a single unified model for heterogeneous multi-dataset learning rather than transfer to a fixed target domain, although DG methods remain complementary and could be layered onto Point-MoE as a natural extension.
3 Point-MoE for Multi-Dataset Point Cloud Training
3.1 Large-Scale Multi-Dataset Joint Training
Task definition. Given a 3D scene represented as a point cloud $\mathcal{P}=\{p_{i}\}_{i=1}^{M}$, where each point $p_{i}\in\mathbb{R}^{3}$ denotes XYZ coordinates, semantic segmentation aims to predict a single class label $\hat{y}_{i}\in\mathcal{C}$ for each point $p_{i}$ from a fixed label set $\mathcal{C}=\{c_{1},c_{2},\ldots,c_{m}\}$. Predicted labels for the entire point cloud are compactly denoted as $\mathbf{\hat{y}}$. Similarly, we denote the corresponding ground-truth labels by $\mathbf{y}=[y_{1},y_{2},\ldots,y_{M}],y_{i}\in\mathcal{C},$ where $y_{i}$ is the true semantic label of point $p_{i}$.
@@ 3.1 Large-Scale Multi-Dataset Joint Training
Let a dataset containing multiple point clouds be denoted as $\mathcal{D}=\mathopen{}\mathclose{{\left\{(\mathcal{P}_{t},\mathbf{y}_{t})}}\right\}_{t=1}^{|\mathcal{D}|}$. The complete training set across multiple datasets is then $\mathbb{D}=\mathopen{}\mathclose{{\left\{\mathcal{D}_{j}}}\right\}_{j=1}^{|\mathbb{D}|}$. The goal of our multi-dataset joint training is to learn a unified model (i.e. a single set of model parameters, without resorting to per-dataset fine-tuning) over all datasets that minimizes prediction error across $\mathbb{D}$, while also generalizing effectively to zero-shot datasets drawn from distributions unseen at training time.
[…]
We can formally contrast this to alternate multi-dataset objective formulations such as [82, 75], that learn an explicit set of dataset-specific parameters $\omega_{j}$, along with $\theta$: $\min_{\theta,\{\omega_{j}\}}\;\mathbb{E}_{\mathcal{D}_{j}\sim\mathbb{D}}\;\mathbb{E}_{(\mathcal{P},\,\mathbf{y})\sim\mathcal{D}_{j}}\mathopen{}\mathclose{{\left[\mathcal{L}\mathopen{}\mathclose{{\left(\mathbf{y},\;\Phi_{\theta,\omega_{j}}(\mathcal{P})}}\right)}}\right]$. PPT [82] instantiates $\{\omega_{j}\}_{j=1}^{|\mathbb{D}|}$ as the learnable gain and bias terms of the normalization layers of PTv3, trained on a per-dataset basis. During inference, the source dataset $\mathcal{D}_{j}$ of an incoming point cloud $\mathcal{P}$ must be identified in order to apply the correct normalization $\omega_{j}$, usually obtained from a lightweight dataset classifier as in One-for-All [75].
@@
Language-guided classification. To bridge label discrepancies across datasets (i.e. $\mathcal{C}_{i}\neq\mathcal{C}_{j}$ for datasets $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$), we project per-point features to a shared language space using CLIP [57] text embeddings, following PPT [82]. This enables supervision via class names, without relying on dataset labels. For instance, the class “pillow” is missing in ScanNet (grouped as “other”) but explicitly labeled in Structured3D; aligning to the CLIP embedding of “pillow” transfers knowledge across datasets even without direct supervision. All baselines and our Point-MoE are adapted to this language-guided protocol, ensuring fair comparison.
3.2 Point-MoE
Mixture-of-Experts (MoE) layer. Mixture-of-Experts [30] is a conditional computation architecture enabling significant capacity expansion while only modestly increasing computational cost. The core of MoE is to route input tokens dynamically to specialized subnetworks, referred to as experts, via a lightweight gating mechanism. Given an input feature vector $\mathbf{x}\in\mathbb{R}^{D}$, each MoE layer contains $N$ experts $\{f_{i}\}_{i=1}^{N}$ where $f_{i}:\mathbb{R}^{D}\!\to\!\mathbb{R}^{D}$ and a gating network $G:\mathbb{R}^{D}\rightarrow\mathbb{R}^{N}$. For each input $\mathbf{x}$, the MoE layer output is computed as a weighted sum over the top-$k$ activated experts:
$\mathrm{MoE}(\mathbf{x})=\sum_{i\in\mathcal{S}_{\mathbf{x}}}G_{i}(\mathbf{x})\,f_{i}(\mathbf{x}),$ (1)
where $\mathcal{S}_{\mathbf{x}}$ is the top-$k$ expert set and $G_{i}(\mathbf{x})$ are normalized gate weights. In practice, $G$ is a linear projection followed by a sparse softmax, activating only $k\!\ll\!N$ experts. Importantly, MoE introduces a large design space, including the choice of which components to replace, the number of experts, the routing strategy, whether to do auxiliary balancing loss and whether to use top-1 or top-2 gating. We systematically study how different MoE configurations affect performance on 3D point clouds through detailed ablations in the experiments.
Integration into PTv3. Given a point cloud $\mathcal{P}$ with $M$ points, PTv3 [81] serializes and embeds the input into a token sequence $\mathbf{Z}=[z_{1},\ldots,z_{M}]$. PTv3 then applies stacked multi-head self-attention (MHSA) blocks interleaved with up- and down-sampling for multi-scale reasoning on these per-token embeddings. At any intermediate layer $\ell$, let $\mathbf{x}_{\ell}\!\in\!\mathbb{R}^{M\times D}$ denote the token sequence at layer $\ell$, where $D$ is the feature dimension, $H$ is the number of heads in multi-head attention, and $\mathrm{FFN}$ denotes a feed-forward network. The computations of a self-attention block using standard QKV-formulation and pre-norm residuals is as follows:
$\mathbf{X}$ $=\mathrm{Norm}(\mathbf{x}_{\ell-1}),\qquad\mathbf{Q}=\mathbf{X}W_{q},\;\mathbf{K}=\mathbf{X}W_{k},\;\mathbf{V}=\mathbf{X}W_{v},$ (2)
$\mathbf{A}$ $=\mathrm{Softmax}\!\big(\mathbf{Q}\mathbf{K}^{\top}/\sqrt{d_{h}}\big)\,\mathbf{V},\quad d_{h}=D/H,$ (3)
$\mathbf{O}$ $=W_{o}\mathbf{A},$ (4)
$\mathbf{x}^{\prime}_{\ell}$ $=\mathbf{x}_{\ell-1}+\mathbf{O},$ (5)
$\mathbf{x}_{\ell}$ $=\mathbf{x}^{\prime}_{\ell}+\mathrm{FFN}\!\big(\mathrm{Norm}(\mathbf{x}^{\prime}_{\ell})\big)$ (6)
The learnable query, key, and value projections are denoted as $W_{q}$, $W_{k}$ and $W_{v}$, while $W_{o}$ denotes the projection applied on the attention output $\mathbf{A}$. We apply MoE only at the attention output projection ($W_{o}$) in every block, rewriting Eq. (4) as $\mathbf{O}=\mathrm{MoE}(\mathbf{A})$ ; the query, key, and value projections ($W_{q},W_{k},W_{v}$) remain dense. Empirically, placing MoE at $W_{o}$ yields better performance than placing it in the FFN under the same expansion (Tab. 3). We hypothesize that the gain comes from adding nonlinearity and capacity on the projection path: with purely linear projections, the composition of $W_{v}$ and $W_{o}$ can be merged into a single projection [47].
Beyond MoE-layer placement, we employ mixed-dataset training: each minibatch jointly samples from both indoor and outdoor data (Tab. 4). Having tokens from multiple datasets within a single minibatch, and thus contributing to the gradients of one update step, is useful in facilitating cross-sample interaction and emergent expert specialization.
To illustrate the general applicability of this framework, we further extend MoE to OA-CNNs [52], a sparse convolution backbone. MoE-augmented OA-CNNs likewise outperform the dense version, further underscoring the effectiveness and general applicability of MoE as it is not tied to a specific backbone architecture. Detailed results are in the supplementary material under Sec. B.2.